OpenStudy (anonymous):

Determine the coordinates of the x-intercepts for the quadratic equation 5x^2 – 21x + 4 = 0.

OpenStudy (anonymous):

Just make the factors of it and then find x..

OpenStudy (anonymous):

i dont know how to factor

OpenStudy (anonymous):

first you multiply the 5 with 4, so you get 20 and then you have to find two numbers that give you 20 when you multiply them and give you a sum of -21 when you add them

OpenStudy (anonymous):

I'll show you an easy example so that you learn how to factorize. \[2x ^{2}+5x+3\]

OpenStudy (anonymous):

2x3=6 and 2+3=5

OpenStudy (anonymous):

so you basically split the middle term into 2x+3x

OpenStudy (anonymous):

now you have a new equation that looks like this \[2x ^{2}+2x+3x+3\]

OpenStudy (anonymous):

then you factorize like this: \[2(x+1)+3(x+1) so the (x+1) is common for both of them

OpenStudy (anonymous):

i forgot the x after the 2

OpenStudy (anonymous):

so it's (2x+3)(x+1)

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

you can check if it's right by just expanding the brackets... do you know how to expand?

OpenStudy (anonymous):

i still dont get it my teacher is kinda teaching me slowly

OpenStudy (anonymous):

okay factorising is when you try to break down the equation to its components

OpenStudy (anonymous):

you do that by finding common factors

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you get this?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

so you need to split the middle term(in my example it's the 5x) to split it means that you have to first multiply the number in front of the x^2 with the last number and then you find a number that will give you the 5 in the middle when you add them and the number you got when you multiplied

OpenStudy (anonymous):

like my example it was 2 and 3... because 2+3=5

OpenStudy (anonymous):

and 2*3=6

OpenStudy (anonymous):

with me?

OpenStudy (anonymous):

yea im with you

OpenStudy (anonymous):

do you get how we split the middle term

OpenStudy (anonymous):

yeah it mean you first muilitply the number in front of x^2

OpenStudy (anonymous):

with the last number which is 3 here

OpenStudy (anonymous):

so you get 6

OpenStudy (anonymous):

you example or my example?

OpenStudy (anonymous):

well my example is easier to understand if you haven't done this before

OpenStudy (anonymous):

and it's much easier if your teacher explains it to you right in front of you so you can see it step by step

OpenStudy (anonymous):

yeah i know, can we use my example?